OpenStudy (anonymous):

In terms of n, find the sum of the first n terms of the arithmetic sequence: -4, 5, 14, 23, 32, .....

OpenStudy (anonymous):

\[a _{n} = a _{1} + (n - 1)d\] \[a _{n} = -4 + 9n - 9 = 9n - 13\] \[S_{n} = n/2(a _{1} + a _{n})\] \[S _{n} = n/2(-4 + 9n - 13)\] \[n(9n - 17)\over2\] \[9n ^{2} - 17n \over2\]